OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = 3x + 1 A. (6, 19) and (0, 1) B. (6, 19) and (0, 0) C. (6, 18) and (0, 1) D. (19, 6) and (0, 1)

OpenStudy (anonymous):

I think its A @AnswerMyQuestions @Love_Ranaa

OpenStudy (igreen):

You are correct.

OpenStudy (anonymous):

plug in the numbers X for 6 and multiply it times 3 and plug in 19 for y and add up and do same thing for the other.....

OpenStudy (anonymous):

Thanks '0'

OpenStudy (igreen):

y = 3x + 1 (6, 19) 19 = 3(6) + 1 19 = 18 + 1 19 = 19 (0, 1) 1 = 3(0) + 1 1 = 0 + 1 1 = 1

OpenStudy (igreen):

\(\Huge\color{lime}\checkmark\)

OpenStudy (love_ranaa):

Ayee..She got it :)

OpenStudy (anonymous):

Thank you ^.^